OpenStudy (anonymous):

how do you find the domain and range of: f(x)= 1/2|x-2|

OpenStudy (mathteacher1729):

Domain = all values of x you can substitute into the equation such that you don't have: 1) Log of a number <= 0 2) Square root of a negative number 3) 1 / 0 Range = All values of y the function reaches when you input all the values of x it can take on.

OpenStudy (anonymous):

I thought it would be (-infinity, 0), things like that

OpenStudy (amistre64):

is this the equation? \[\frac{1}{2|x-2|}\]

OpenStudy (anonymous):

no

OpenStudy (anonymous):

it's the way I posted it

OpenStudy (amistre64):

:) the way you posted it can be interpreted a number of different ways

OpenStudy (amistre64):

\[\frac{1}{2}|x-2|\] if this is it you should have posted it like this: (1/2) |x-2|

OpenStudy (anonymous):

I was told by my teacher that the domain is from (-infinity,infinity) I don't even know the range. I'm so lost as to why they gt these as answers. Especially since the grapgh of an | | equation looks like this

OpenStudy (amistre64):

the range is the lowest value up to the highest value

OpenStudy (anonymous):

Yes, the equation is: 1/2 times the absolute valuse ||, of x-2- |x-2|

OpenStudy (amistre64):

and that is not the graph of what you have posted; your equation has a horizontal shift, not a vertical shift

OpenStudy (anonymous):

I know, I was just saying that an absolute value graph looks like that

OpenStudy (mathteacher1729):

The domain of |x| is all values of x. The RANGE of |x| is all values of x greater than or equal to zero. (obviously, you can't get a negative y-value by taking the absolute value of x. 1/2|x| is just "stretching" the graph, domain and range remain unchanged. 1/2|x-2| is shifting the vertex two units to the right. (Because x = 2 gives y = 0 ). The domain and range remain unchanged. Domain = all real numbers Range = all real numbers GREATER THAN OR EQUAL TO zero. :)

OpenStudy (amistre64):

that would be one drawing yes :) But the equation you have just moves side to side; so the range is unaffected; it remains (0,inf)

OpenStudy (mathteacher1729):

The blue function is 1/(2|x-2|), which has an asymptote at x = 2 and the domain is all real numbers EXCEPT x = 2.

myininaya (myininaya):

thats pretty teacher

OpenStudy (mathteacher1729):

Thanks, I made it in geogebra, I'm giving a demo of it now if you or anyone else is interested. :)

OpenStudy (anonymous):

Okay, so based on your gorgeous graph, here are my questions.

OpenStudy (anonymous):

Wouldn't one of the domins be: (-infinity,0)

OpenStudy (anonymous):

An another (0,2)

myininaya (myininaya):

when you find the domain look from left to right the domain is where the function exist

myininaya (myininaya):

is there any real input that you can think of that would not give you a real output?

OpenStudy (anonymous):

well I know that it'd an infinite graph

myininaya (myininaya):

the domain is all real numbers because for each of these real numbers there is a real output (-inf,inf)=the set of all real numbers

myininaya (myininaya):

if we had 1/(2|x-2|), the domain would be all real numbers except x=2 since that would give us zero on the bottom

OpenStudy (mathteacher1729):

Myininaya -- great explanations. :)

myininaya (myininaya):

thanks :)